[(z^2 - 25) \ z] \ [(z - 5) \ (z - 8)]
Answer by jorel1380(3719)

(z2-25)/z divide by (z-5)/(z-8)=
z2-25/z X (z-8)/(z-5)=(z+5)(z-8)/z=z2-3z-40/z..